OpenStudy (anonymous):

i will give a medal if you can help me with this one problem!!!geometry 13.Write an equation in point-slope form of the line throught J(4,1) with slope -4. if you dont want to answer it atleast walk me through it please!!:)

OpenStudy (anonymous):

I'll walk you through it. So what does the point-slope form look like?

OpenStudy (anonymous):

all i know is the given information -Write an equation in point-slope form of the line throught J(4,1) with slope -4. i dont know how to write the equation in point-slope form

OpenStudy (anonymous):

so point- slope form is given by this formula: \[y-y_{1} = m(x - x_{1})\] where m is the slope and x1,y1 is a point on the line. So given that information can you write the solution to your problem?

OpenStudy (anonymous):

would it be y-4=-4(x-1)?

OpenStudy (anonymous):

close! you switched the x and y coordinates!

OpenStudy (anonymous):

so y-=-4(x-4)?

OpenStudy (anonymous):

y-1=-4(x-4)?***

OpenStudy (anonymous):

exactly good job!

OpenStudy (anonymous):

thank you so much!!:)

OpenStudy (anonymous):

no problem